OpenStudy (anonymous):

Find the equation of the inverse of each given function. F(x): (x+1)^3 I messed up in my calculating but i cant find the error

OpenStudy (anonymous):

I did y=(x+1)^3 Then Y=x^3 +1^3 Subtract 1^3 Then i got y-1=x ^3 but then i got messed up from there

OpenStudy (anonymous):

Actually wait should i have left -1 on that side ok never mind got it but wait i have a question

OpenStudy (imstuck):

\[y=(x+1)^{3}\]

OpenStudy (imstuck):

replace x and y with each other and solve for y.

OpenStudy (imstuck):

\[x = (y+1)^{3}\]

OpenStudy (anonymous):

And then solve to isolate y

OpenStudy (imstuck):

to undo the cubing of the (y+1) you have to take the cubed root of both sides, like this:\[\sqrt[3]{x}=y+1\]Now subtract the 1 from both sides so you have y alone:\[\sqrt[3]{x}-1=y\]

OpenStudy (anonymous):

Yea i got that already haha, just had to think it through... I have a massive headache because of school

OpenStudy (anonymous):

But when want to test an equation for symmetry algebraically would you just negate the value of f(x) so it would be f(-x) ? And how do you determine where it reflects ?